OpenStudy (breanneh):

Help ! I'll fan and medal ! Which ordered pair is NOT a solution of y=-12x-3 ?? A. (-3,33) B. (8,-98) . C. (-9,105) D. (3,-39)

jimthompson5910 (jim_thompson5910):

Plug in each ordered pair The ordered pair that gives you a false equation is the answer you want

jimthompson5910 (jim_thompson5910):

For example Choice A (-3,33) means x = -3 and y = 33 Plug them into y=-12x-3 to get y=-12x-3 33=-12(-3)-3 33= 36 - 3 33 = 33 ... this equation is true so (-3,33) is a solution That rules out choice A because we want something that's not a solution

OpenStudy (anonymous):

The above method is right , what do you want to know?